Write a Pythagorean triplet whose one member is. 16

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

For any natural number m, we know that 2m, m2–1, m2+1 is a Pythagorean triplet.
2m=16⇒m=16/2=8
m^2–1=8^2–1=64–1=63
m^2+1=8^2+1=64+1=65
∴ (16, 63, 65) is aPythagorean triplet.

​Thanks